OpenStudy (anonymous):

@math&ing001 Sorry to put you to work xD If two squares have the same area,what do you know about the measures of their sides?Explain.

OpenStudy (math&ing001):

They would have equal sides. Think of it as an equation: first square with side=x Second square with side=y x*x = y*y => x=y (x and y >0)

OpenStudy (anonymous):

Okaiii so what if two rectangles have the same area,What do you know about the measures of their sides??

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

The product of their sides are equal ab=cd 

OpenStudy (anonymous):

@math&ing001 That makes Perfect since,i was going to say something like they were adjacent to one another.